how many gallons of a 60% antifreeze solution must be mixed with 90 gallons of 30% antifreeze to get a mixture that is 50% antifreeze?
Answer by stanbon(75887)

how many gallons of a 60% antifreeze solution must be mixed with 90 gallons of 30% antifreeze to get a mixture that is 50% antifreeze?
------
Equation:
anti + anti = anti
0.30*90 + 0.60x = 0.50(90+x)
------
Multiply thru by 100 to get:
30*90 + 60x = 50*90 + 50x
10x = 20*90
x = 180 gallons (amt. of 60% mix to add)
=========================
Cheers,
Stan H.